# Development first Performance evaluation of multi element monolithic HPGe detector for spectroscopy applications

### v1

@@ 3.3.1 Energy Resolution Count Rate Effects in Fluorescence Mode
We used samples of Mn, Sn, and BaO, producing characteristic X-ray lines ranging from 6 to 33 keV. At low input count rates (10 kcps), the detector achieved optimal resolution at shaping times between 4 and 6 µs, where the noise contribution was minimal, and charge collection was complete. As the ICR increased to moderate levels (200–300 kcps), the optimal shaping time shifted slightly toward 2–4 µs, balancing resolution and event throughput. At high ICRs (1.2 Mcps), resolution degraded significantly for PT values above 6 µs, primarily due to higher dead time, while shorter PTs provided better resolution in a higher flux environment, as shown in Fig 8.
@@ 3.3.2 Dead time and pile-up studies
We characterized the detector’s performance under high flux conditions by building a relationship between the input count rate (ICR) and the output count rate (OCR). Since incoming photons follow a Poisson process, the time interval between successive arrivals is exponentially distributed. For a spectroscopic detector, an event must be sufficiently isolated in time to be processed as an individual pulse; if another photon arrives too close in time, the two pulses sum together (pile-up), hence distorting the shape of the energy spectrum [blaj2017optimal].
@@
where is the detector’s effective processing time. This expression follows directly from the Poisson probability of observing no additional photons within the processing window of duration 2. The parameter embeds the properties of the slow channel of the pulse processor. This simple model was used to fit from our measured OCR and ICR data. Fitted values are in good agreement with the setting of the pulse processor.
The time constant of the pulse processor fast channel can also be extracted from the data. At high ICR, pile-up peaks appear at twice the energy of the main fluorescence line, i.e., the fast channel is unable to resolve two or more photons arriving as a bunch within a short time interval, as seen in Fig. 9(a). According to Poisson statistics, the output count rate of pileup events involving 2 photons is given by the expression . By estimating the output count rate ratio between the pileup events involving 2 photons and the count rate of single events, we could estimate the time constant of the pulse processor fast channel: electronics estimated the ratio of the fitted the main peak and the corresponding pile-up peak with Gaussian functions, and defined the pile-up ratio as
@@
From our measurements, we obtain a consistent value of , see Fig 9(b). Extended models which includes additional parameters such as a reset duration term of CSA have been proposed in the literature [Bordessoule2019]. These models account for the effect of the preamplifier reset dead time at very high flux or high photon energy. In our measurements, no significant saturation behaviour was observed. The data remained in a regime where the response can be well described by using a simple paralyzable model, without adding any additional saturation parameter, as shown in Fig 9(c). We initially tested models that included the CSA reset duration; however, the data are well described by a simple paralyzable model, indicating that the CSA reset duration had a negligible effect on the detector’s OCR in the investigated regime.
@@ 3.4 Surface scanning: Micro spot beam
To investigate the spatial response of the XAFS-DET prototype, we performed surface scans using the microspot X-ray beam from the generator installed at the ESRF’s detector laboratory. The experimental setup, schematically illustrated in Figure 10, consisted of a copper anode X-ray source delivering an 8.04 keV beam. The beam was collimated using motorized slits, followed by a titanium pinhole of 100 m diameter to achieve a well-defined spot on the detector surface. Beam flux was tuned by inserting copper filters in the beam path.
The detector was mounted on a motorized XZ stage for automated scanning and was biased at +70 V throughout the measurements. Four pixels were read out simultaneously via an XIA-Mercury digital pulse processor (DPP), allowing us to use the TTL outputs from the user-defined region of interest (ROI) in the energy spectrum. It enabled easy integration with external counting systems. The DPP was configured with a single-channel analyzer (SCA) to count only events in the Cu K peak ROI at 8.04 keV, see Fig. 10. Surface mapping was performed with a step size of 50 m and an exposure time of 1 s per position, covering three to four pixels per scan area. The main scan parameters are summarized in Table 2.
@@
| Number of scans | 4-5 (covering 3–4 pixels per scan) |
The reconstructed surface map obtained by combining individual scans is shown in Figure 11(a). To create a spatial response map of the entire sensor, data from multiple, smaller raster scans were combined into a single heatmap. Since individual scans covered only a few pixels or a specific region of the detector, this "patchwork" approach was essential for visualizing the performance across the full sensor surface. The reconstruction process was first performed by compiling all the individual scans, along with the motor coordinates (hy, hz) and the corresponding signal counts for each active detector pixel. We then aggregated this information into a single dataset to create a 2-D histogram, which served as the canvas for the final image. A uniform bin width of 50 µm was chosen to match the high-resolution scan step size. The ranges of the histogram’s x and y axes were set to cover the full extent of all motor positions recorded across all scans. The aggregated data points were then binned into the 2D histogram. The value of each bin in the final map was determined by the total number of counts recorded at that specific (x, y) spatial coordinate, summed across all contributing pixels from all scans. As a result, we confirmed that the pixel boundaries closely match the molybdenum collimator apertures, confirming excellent alignment between the collimator and the sensor.
Four distinct regions of reduced sensitivity were detected (highlighted in red), see Fig 11(a), corresponding to the defective area in the germanium sensor. When the X-ray beam (count rates of 30-35 kcps) illuminated these areas, the leakage current increased from its nominal value of 1 pA to approximately 100 pA.
@@
To further investigate the defective areas, energy spectra were acquired at positions within (black regions) and outside (yellow region) each defect. A typical example is shown in Figure 11(b). The spectrum from within the defect shows reduced total counts and a suppressed Cu K peak, whereas the normal region shows the expected photopeak with nominal intensity. These results confirm that the observed low-response zones in the surface map correspond to true inactive regions in the sensor.
@@ 3.4.1 Crosstalk between pixels
To study the crosstalk between pixels, we positioned the beam to cover the interpixel spacing, i.e., the collimator edge. In this position, three to four pixels were directly illuminated, while the rest of the neighboring pixels received no direct X-ray beam. In this measurement, for each scan, one non-illuminated pixel was connected to the readout electronics if the scanned area covered no more than three pixels. Crosstalk analysis, illustrated in Figure 12(a), revealed that when one pixel was scanned, a neighboring pixel not directly illuminated registered a faint shadow of the collimator.
@@ 3.4.2 Rise Time Scanning
We evaluated the timing performance by performing Rise-time scans using a digital oscilloscope coupled to the detector output. The oscilloscope was configured with a horizontal time base of 200 ns/div and a vertical sensitivity of 20–40 mV/div (depending on signal amplitude). Data were acquired as fall-time histograms, with each mesh step corresponding to a 4 s integration time. The spatial scan was performed in a 2D mesh pattern, using a 50 µm step size. We measured the signal risetime distribution for three pixels, as shown in Figure 13. The average risetime was approximately 80 ns at +70 V, consistent across the pixel surface, except for some edge effects, visible along the edges of three pixels. The measured risetime is consistent with the simulated drift time of charge carriers in the sensor (50-70 ns at +70V).
@@ 4 First On-beam tests
The first synchrotron measurements with our 1st prototype detector were performed at the BM05 beamline in the ESRF-EBS facility [BM05beamline]. The BM05 beamline provides an experimental hutch with a focused monochromatic X-ray beam, achieving a spot size of approximately 35 µm for detector characterization and follow-up spectroscopy tests.
The detector was mounted on a two-axis motorized stage for precise alignment with the incident beam. In the same hutch, a dedicated diagnostics stage was equipped with a standard photodiode for monitoring beam intensity and a Basler camera to visualize the beam spot size. A schematic representation of the final setup is shown in Fig. 14.
@@
| Beam Spot Size | ~35 µm |
@@
| Low-Resolution Scan Step | 250 µm |
| High-Resolution Scan Step | 25 µm |
@@ 4.1 Performance at Monochromatic Energies
During the beam tests, the detector was exposed to a focused, monochromatic X-ray beam. We acquired several two-dimensional spatial response maps as well as the energy spectrum for each scan point to assess detector performance and investigate the localized defects previously identified in laboratory microsource beam scans. We performed high-resolution X-ray scans by exposing the sensor surface, and specifically the defect area, to high-energy photons and by exploring their behavior under different detector biases. The idea of modeling the suppression to estimate defect depth emerged later, followed by the curiosity about the origin of the defects. This led to the development of a simple yet effective model that enabled us to estimate the depth of the defects from the suppression data of each scan. The beam and the scan parameters are tabulated in Tab. 3
Two types of scans were performed. First, low-resolution raster scans (250 µm step size) across the entire sensor area at the three aforementioned energies to generate a comprehensive performance map. These scans showed several defective regions, as evidenced by pixels 1, 5, 8, and 9, which exhibited significant signal loss. Following this, we acquired high-resolution scans (25 µm step size), focusing mainly on the defective areas of pixel 5 and 9. These targeted scans were conducted at three bias voltages to obtain the necessary data for a quantitative analysis of the defect’s behaviour.
@@ 4.2 Defect depth estimation
With this analysis, we estimated the physical depth of the defects based on measured signal suppression. We used two-dimensional heatmaps of the defect region to define an ROI containing the defect. The defect severity is evaluated using the signal suppression ratio , defined as the ratio of the average counts in the defective region to the maximum counts in the good region , as shown using eq. 4.1.
The selected ROIs are also highlighted using Fig. 17. In our calculations, one data point was excluded from the final suppression ratio fit
333The 20 keV scan of pixel 9 at 70 V was excluded because there was no clear defect contrast.. For better understanding, we also used an illustration of the photon interaction point, charge transport, and the presence of defects in the Ge material in Fig. 16.
Since both and were measured under the same conditions with the same entrance windows (Be Kapton) in the beam path, the corresponding transmission losses cancelled out, and no explicit correction by a transmission factor is included. The measured suppression ratio reflects the charge collection efficiency between the defect and the normal region of the pixels. To convert this into an effective physical depth , we used a phenomenological model that accounts for both photon attenuation and incomplete charge collection.
The suppression of the X-ray signal in the presence of a defect arises from two physical effects: attenuation of the incident photon as it travels deeper into the material, and a reduction in charge-collection efficiency due to carrier trapping. We model this phenomenon as:
In this expression, is the energy-dependent linear attenuation coefficient of Ge (obtained from xraylib [brunetti2004library]), while represents the mean drift length which accounts for the average distance an electron-hole pair can travel before being trapped. Since a stronger field improves the charge transport and therefore the charge collection, we model as a voltage-dependent parameter, increasing with the applied bias :
In this formulation, the two parameters ( and ) are effective parameters that capture the voltage-dependent charge-collection behavior of the Ge bulk. Physically, represents the intrinsic mean drift length of charge carriers at zero applied bias, dictated by intrinsic material quality such as sensor impurities or the density of trapping centers.
The quantifies how the effective collection depth increases with the electric field, showing a reduction in the trapping probability and an increase in the carrier drift velocity at higher bias. Both parameters were fitted globally across all bias and energy values by minimizing the deviation between the measured and predicted suppression ratios and the model, respectively. The best-fit parameters were found to be mm and V-1. For each (energy, bias) point, the defect depth was obtained by numerically inverting the suppression model given in eq. 4.2, by solving
@@
We summarize this analysis results using four subfigures of Fig. 18.
@@
Figures (b) and (c) show the measured suppression ratios and defect areas, respectively, showing that the defect in pixel 5 is severe as compared to the one observed in pixel 9. With increasing bias, both defects shows improved charge collection, resulting in less suppression and a smaller area.
Finally, Figure (d), shows the model curves from the CCE model using the fitted and , overlaid with measured data points. The shape of the model curves can be interpreted from the competing product of photon attenuation and charge collection efficiency factor as shown in the model expression 4.2. For shallow defects (), the suppression rises linearly with depth since only a small fraction of created charge is lost. As approaches , the charge collection saturates and flattens, showing the transition from partial to nearly complete charge collection. At large depths, the exponential attenuation term dominates, leading to a decrease at higher energies. Increasing the detector bias effectively extends , leading to shallower and less pronounced suppression curves, consistent with improved charge transport in stronger electric fields.
@@
In addition, to verify that the depths extracted with this model are physically consistent with the photon interaction range, we also calculated the photon-absorption distribution in Ge using energy-dependent attenuation coefficients from Xraylib. The calculation also includes transmission losses through the Be and Kapton entrance window and is shown in Table 4. The resultant mean absorption depths (0.03–0.56 mm for 8–50 keV) are orders of magnitude greater than the fitted defect depths (1–6 µm), confirming that the observed suppression originates from shallow regions near the sensor’s entrance window rather than from absorption limits or deep bulk trapping. A summary of used data points and the estimated defect height, width and depth is tabulated using Tab. 5.
@@ 4.3 Dead Time evaluation with Direct beam
During the beam time at the BM05 beamline, we evaluated the detector’s count rate performance in high-flux conditions (2-3 Mcps) by exposing it to a direct monochromatic X-ray beam. Measurements were taken at two shaping time values: 1 µs (targeted for high rate operation) and 6 µs (currently used for optimal energy resolution). Three ASIC gain settings were tested to study their effect on the count rate behavior. The data below correspond to the Pixel 3, which has shown no signs of defects up to an exposure of a 50 keV photon beam.
The input count rate (ICR) was varied by adjusting the beamline slits, while maintaining a constant beam spot size of approximately 35 µm on the sensor. The ICR was derived from a standard photodiode, while the output count rate (OCR) was measured from the digital pulse processor (DPP).
@@
We observe that at a PT value of 6 µsec, the OCR saturates at lower ICR values (30-40 kcps), showing significant dead time losses, whereas, at 1 µs PT, the detector possessed linear response up in a broader range, and began saturating at ICR values above 150-200 kcps The extracted values from the beamtime data agree well with the values obtained from the lab fluorescence measurements in Sec. 3.3.2.

### v2

@@ 3.3.1 Energy Resolution Count Rate Effects in Fluorescence Mode
The sample foils used for characterization were selected to test and cover the three ASIC gain settings. The fluorescence spectra were calibrated using the characteristic X-ray lines at 5.9 keV(Mn-), 25.27 keV (Sn-), and 32.1 keV (Ba-). The corresponding energy resolution are also showcased using Figures 8(a), 8(c), 8(e). At low input count rates (10 kcps), the detector achieved optimal resolution at shaping times between 4 and 6 s, where the noise contribution was minimal, and ballistic deficit was minimal. As the ICR increased to moderate levels (200–300 kcps), the optimal shaping time shifted slightly toward 2–4 s, balancing resolution and event throughput. At high ICRs (1.2 Mcps), resolution degraded significantly for PT values above 6 s, while shorter PTs provided better resolution in a higher flux environment, as shown in Figures 8(b), 8(d), 8(f).
@@ 3.3.2 Dead time and pile-up studies
We characterized the detector’s performance under high-flux conditions by establishing a relationship between the input count rate (ICR) and the output count rate (OCR). Since incoming photons follow a Poisson process, the time interval between successive arrivals is exponentially distributed. For a spectroscopic detector, an event must be sufficiently isolated in time to be processed as an individual pulse; if another photon arrives too close in time, the two pulses sum together (pile-up), hence distorting the shape of the energy spectrum [blaj2017optimal].
@@
where is the detector’s effective processing time. This expression follows directly from the Poisson probability of observing no additional photons within the processing window of duration 2. The parameter embeds the properties of the slow channel of the pulse processor. This simple model was used to fit from our measured OCR and ICR data. Fitted values are in good agreement with the pulse processor settings.
The time constant of the pulse processor fast channel can also be extracted from the data. At high ICR, pile-up peaks appear at twice the energy of the main fluorescence line, i.e., the fast channel is unable to resolve two or more photons arriving as a bunch within a short time interval, as seen in Figure 9(a). According to Poisson statistics, the output count rate of pileup events involving 2 photons is given by the expression :
By estimating the output count rate ratio between the pileup events involving 2 photons and the count rate of single events, we could estimate the time constant of the pulse processor fast channel: electronics estimated the ratio of the fitted the main peak and the corresponding pile-up peak with Gaussian functions, and defined the pile-up ratio as
@@
From our measurements, we obtained a for all PTs, see Figure 9(b). Extended models that include additional parameters such as a reset duration term of CSA have been proposed in the literature [Bordessoule2019]. These models account for the effect of the preamplifier reset dead time at very high flux or higher photon energy. In our measurements, no significant saturation behaviour was observed. The data remained in a regime in which the response could be well described by a simple paralyzable model, without requiring an additional saturation parameter, as shown in Figure 9(c). We initially tested models that included the CSA reset duration. however, the data are well described by a simple paralyzable model, indicating that the CSA reset duration had a negligible effect on the detector’s OCR in the investigated regime.
@@ 3.4 Surface scanning: Micro spot beam
To investigate the spatial response, we performed surface scans using the microspot X-ray beam from the generator installed at the ESRF’s detector unit laboratory. The experimental setup, schematically illustrated in Figure 10, consisted of a copper anode X-ray source delivering an 8.04 keV beam. The beam was collimated using motorized slits, followed by a 100 m-diameter titanium pinhole to achieve a well-defined spot on the detector surface. Beam flux was tuned by inserting copper filters of varying thickness in the beam path.
The detector was mounted on a motorized X-Z stage for automated scanning and was biased at +70 V throughout the measurements. Four pixels were read out simultaneously via an XIA-Mercury digital pulse processor (DPP), allowing us to use the TTL outputs from the user-defined region of interest (ROI) in the energy spectrum. It enabled easy integration with external counting systems which count only events within the Cu K peak ROI at 8.04 keV, see Figure 10. We performed the mapping of sensor surface with a scan step size of 50 m and an exposure time of 1 s per position, covering three to four pixels per scan area. The main scan parameters are summarized using Table 2.
@@
| Number of scans | 4-5 (covering 3–4 pixels/scan) |
The reconstructed surface map obtained by combining individual scans is shown in Figure 11(a). To create a spatial response map of the entire sensor, data from multiple, smaller raster scans were combined into a single heatmap. Since individual scans covered only a few pixels or a specific region of the detector, this "patchwork" approach was essential for visualizing the performance across the full sensor surface. The reconstruction process was first performed by compiling all the individual scans, along with the motor coordinates (hy, hz) and the corresponding signal counts for each active detector pixel. We then aggregated this information into a single dataset to create a 2-D histogram, which served as the canvas for the final image. A uniform bin width of 50 m was chosen to match the high-resolution scan step size. The ranges of the histogram’s x and y axes were set to cover the full extent of all motor positions recorded across all scans. The aggregated data points were then binned into the 2D histogram. The value of each bin in the final map was determined by the total number of counts recorded at that specific (x, y) spatial coordinate, summed across all contributing pixels from all scans. As a result, we confirmed that the pixel boundaries closely match the molybdenum collimator apertures, confirming excellent alignment between the collimator and the sensor.
Four distinct regions of reduced sensitivity were detected (highlighted in red), see Figure 11(a), corresponding to the defective area in the germanium sensor. When the X-ray beam (count rates of 30-35 kcps) illuminated these areas, the leakage current increased from its nominal value of 1 pA to approximately 100 pA.
@@
To further investigate the defective areas, energy spectra were acquired at positions within (black regions) and outside (yellow region) each defect. A typical example is shown in Figure 11(b). The spectrum from within the defect shows reduced total counts and a suppressed Cu-Kα peak, whereas the normal region shows the expected photopeak with nominal intensity. These results confirm that the observed low-response zones in the surface map correspond to true inactive regions in the sensor.
@@ 3.4.1 Crosstalk between pixels
To study the crosstalk between pixels, we positioned the beam to cover the interpixel spacing, i.e., the collimator edge. In this position, three out of four pixels (max.) were directly illuminated, while the rest of the neighboring pixels received no direct exposure. In this measurement, for each scan, one non-illuminated pixel was connected to the readout electronics. Crosstalk analysis, illustrated in Figure 12(a), revealed that when one pixel was scanned, a neighboring pixel not directly illuminated registered a faint shadow of the collimator.
@@ 3.4.2 Rise Time Scanning
We evaluated the timing performance by performing Rise-time scans using a digital oscilloscope coupled to the detector output. The oscilloscope was configured with a horizontal time base of 200 ns/div and a vertical sensitivity of 20–40 mV/div (depending on signal amplitude). Data were acquired as fall-time histograms, with each mesh step corresponding to a 4 s integration time. The spatial scan was performed in a 2D mesh pattern, using a 50 m step size. We measured the signal risetime distribution for three pixels, as shown in Figure 13. The average risetime was approximately 80 ns at +70 V, consistent across the pixel surface, except for some edge effects, visible along the edges of three pixels. The measured risetime is consistent with the simulated drift time of charge carriers in the sensor (50-70 ns at +70V).
@@ 4 First On-beam tests
The first synchrotron measurements with our 1st prototype detector were performed at the BM05 beamline in the ESRF-EBS facility [BM05beamline]. The BM05 beamline provides an experimental hutch with a focused monochromatic X-ray beam, achieving a spot size of approximately 35 m for detector characterization and follow-up spectroscopy tests.
The detector was mounted on a two-axis motorized stage for precise alignment with the incident beam. In the same hutch, a dedicated diagnostics stage was equipped with a standard photodiode for monitoring beam intensity and a Basler camera to visualize the beam spot size. A schematic illustration of the final setup is presented in Figure 14.
@@
| Beam Spot Size | ~35 m |
@@
| Low-Resolution Scan Step | 250 m |
| High-Resolution Scan Step | 25 m |
@@ 4.1 Performance at Monochromatic Energies
During the first beam tests, we acquired several two-dimensional spatial response maps as well as the energy spectrum, (see Figure 15(a)), for each scan point to assess detector performance (energy resolution at 40 keV, shown in Figure 15(c)), and investigate the localized defects previously identified in laboratory micro source beam scans. We performed high-resolution X-ray scans by exposing the sensor surface, and specifically the defect area, to high-energy photons and by exploring their behavior under different detector biases. The idea of modeling the suppression to estimate defect depth emerged later, followed by the curiosity about the origin of the defects. This led to the development of a simple yet effective model that enabled us to estimate the depth of the defects from the suppression data of each scan. The beam and the scan parameters are tabulated in Table 3.
Two types of scans were performed. First, low-resolution raster scans (250 m step size) across the entire sensor area at the three aforementioned energies to generate a comprehensive performance map. These scans showed several defective regions, as evidenced by pixels 1, 5, 8, and 9, which exhibited significant signal loss, as shown in Figure 15(b). Following this, we acquired high-resolution scans (25 m step size), focusing mainly on the defective areas of pixel 5 and 9. These targeted scans were conducted at three bias voltages to obtain the necessary data for a quantitative analysis of the defect’s behaviour, as shown in Figure 16.
@@ 4.2 Defect depth estimation
With this analysis, we estimated the physical depth of the defects based on measured signal suppression. We used two-dimensional heatmaps of the defect region to define an ROI containing the defect. The severity of the defect is evaluated using the signal/count suppression ratio , defined as the ratio of the average counts in the defective region to the maximum counts in the good region , as shown using eq. 4.1.
The suppression of the X-ray signal in the presence of a defect arises from two physical effects: absorption of the incident photon as it travels deeper into the material, and a reduction in charge-collection efficiency due to carrier trapping. We model this phenomenon as:
In this expression, is the energy-dependent linear attenuation coefficient of germanium (obtained from xraylib [brunetti2004library]), while represents the mean drift length, which accounts for the average distance an electron-hole pair can travel before being trapped. Since a stronger field improves the charge transport and therefore the charge collection, we model as a voltage-dependent parameter, increasing with the applied bias :
In this formulation, the two parameters ( and ) are effective parameters that capture the voltage-dependent charge-collection behavior of the germanium bulk. Physically, represents the intrinsic mean drift length of charge carriers at zero applied bias, dictated by intrinsic material quality such as sensor impurities or the density of trapping centers.
The quantifies how the effective collection depth increases with the electric field, showing a reduction in the trapping probability and an increase in the carrier drift velocity at higher bias. Both parameters were fitted globally across all bias and energy values by minimizing the deviation between the measured and predicted suppression ratios and the model, respectively. A similar approach has been shown in The best-fit parameters were found to be mm and V-1. For each (energy, bias) point, the defect depth was obtained by numerically inverting the suppression model given in eq. 4.2, by solving
@@
We summarize these analysis results using four subfigures of Figure 17.
@@
Figures (b) and (c) show the measured suppression ratios and defect areas, respectively, showing that the defect in pixel 5 is severe as compared to the one observed in pixel 9. With increasing bias, both defects shows improved charge collection, resulting in less suppression and a smaller area. In our calculations, one data point was excluded from the final suppression ratio fit
222The 20 keV scan of pixel 9 at 70 V was excluded because there was no clear defect contrast..
Finally, Figure (d), shows the model curves from the CCE model using the fitted and , overlaid with measured data points. The shape of the model curves can be interpreted from the competing product of photon attenuation and charge collection efficiency factor as shown in the model expression 4.2. For shallow defects (), the suppression rises linearly with depth since only a small fraction of created charge is lost. As approaches , the charge collection saturates and flattens, showing the transition from partial to nearly complete charge collection.
At large depths, the exponential attenuation term dominates, leading to a decrease at higher energies.
Increasing the detector bias effectively extends , leading to shallower and less pronounced suppression curves, consistent with improved charge transport in stronger electric fields.
@@
In addition, to verify that the depths extracted with this model are physically consistent with the photon interaction range, we also calculated the photon-absorption distribution in germanium using energy-dependent attenuation coefficients from Xraylib. The calculation also includes transmission losses through the Be and Kapton entrance window and is shown in Table 4. The resulting mean absorption depths (0.03–0.56 mm for 8–50 keV) are orders of magnitude greater than the fitted defect depths (1–6 m), confirming that the observed suppression originates from shallow regions near the sensor’s entrance window rather than from absorption limits or deep bulk trapping. A summary of used data points and the estimated defect height, width, and depth is tabulated using Table 5.
@@ 4.3 Dead Time evaluation with Direct beam
During the beam time at the BM05 beamline, we evaluated the detector’s count rate performance in high-flux conditions (2-3 Mcps) by exposing it to a direct monochromatic X-ray beam. Measurements were taken at two shaping time values: 1 s (targeted for high rate operation) and 6 s (currently used for optimal energy resolution). Three ASIC gain settings were tested to study their effect on the count rate behavior. The data below correspond to the pixel 3, which has shown no signs of defects up to an exposure of a 50 keV photon beam.
The input count rate (ICR) was varied by adjusting the beamline slits, while maintaining a constant beam spot size of approximately 35 m on the sensor. The ICR was derived from a standard photodiode, while the output count rate (OCR) was measured from the digital pulse processor (DPP).
@@
We observe that at a PT value of 6 s, the OCR saturates at lower ICR values (30-40 kcps), delivering significant dead time losses, whereas, for 1 s PT, the detector maintained linear response up in a broader range, and began saturating at ICR values above 150-200 kcps The extracted values from the beamtime data agree well with the values (2.5 and 13 s) at 1 and 6 s PT, respectively, obtained from the lab fluorescence measurements in Section 3.3.2.